# Z-REx: Human-Interpretable GNN Explanations for Real Estate Recommendations

### v1

@@ 4.3. ZiGNN Architecture
5. Z-REx Overview
The workflow of Z-REx is illustrated in Figure 4 where Z-REx first performs feature perturbation to find the important features and then performs graph structural perturbations to identify the important subgraph structures. We focus on user-city due to the problem definition in § 3, but this method works for any relationship.
5.1. Feature Perturbation.
To interpret the GNN-based recommendation system, we first analyze the influence of target node features through a feature perturbation approach. Consider a user $u$ and their recommended city $c_{t}$ with their embeddings, $\mathbf{h}_{u}$ and $\mathbf{h}_{c_{t}}$. The embeddings are obtained from the trained ZiGNN model and their similarity is computed using cosine similarity:
The feature perturbation involves the following steps:
  1. (1)
  2. (2)
5.2. Structural Perturbation.
While feature perturbation focuses on node-level characteristics, structural perturbation analyzes the impact of graph topology on the model’s predictions. Thus, accounting for the whole graph context into each recommendation explanation. Please note that structural perturbation happens with only the subset of features identified in the previous step. By combining feature and structural perturbation techniques, we provide a comprehensive and interpretable explanation of the GNN-based product recommendation system. Feature perturbation identifies node attributes with significant influence on the recommendation outcomes, while structural perturbation uncovers graph edges and relationships critical to the model’s predictions.
  1. (1)
  2. (2)
    Identify Co-clicked Cities: For the subgraph $\mathcal{G}_{u}^{k}$, we identify pairs of cities $(c_{i},c_{j})$ that share a common predecessor user $u_{p}$. These pairs are added as new edges, representing co-click relationships, and their contributions to the model predictions are evaluated.
  3. (3)
Edges with the highest absolute $\Delta\text{sim}$ values are identified as critical contributors to the recommendation. These edges represent strong graph relationships that drive user preferences for specific cities. The hyperparameters that influence the structural perturbations are: (1) $k$ (hop distance) as it limits the number of edges perturbed to efficiently evaluate while focusing on the most impactful connections and increasing $k$ captures more indirect relationships but may introduce noise, and (2) edge removal strategy as prioritizing edges based on shared predecessors ensures that only influential connections are analyzed. Hyperparameters such as $k$ (neighborhood size) and edge prioritization strategies enable a flexible and robust explanation framework, balancing depth and computational efficiency.
6. Evaluation
6.1. Methodology
The data processing of ZiGNN begins by extracting features containing user activities (e.g., views and saves), user and listing attributes, and their geographical regions. During preprocessing, missing values are handled systematically to ensure data integrity. Outliers, which can distort analysis and model training, are addressed by replacing extreme values (detected using Z-scores) with the column mean. These steps ensure a clean and reliable dataset. Additionally, numerical features are normalized using z-score normalization, creating a consistent data representation. These preprocessing steps create a robust and standardized dataset, ready for graph-based modeling in the recommendation system.
After preprocessing, the user, listing, and city data are used to construct a heterogeneous graph. Each entity (e.g., user, listing, and city) is represented as a unique node type, and edges capture interactions between these nodes. For instance, user-to-listing edges represent activities such as views or saves, while city-to-listing edges denote listings contained within a city. The graph construction also includes bidirectional edges to preserve information flow in both directions, enhancing the ZiGNN’s ability to learn relationships. Node features are assigned and this comprehensive graph representation forms the backbone of the recommendation system, enabling the ZiGNN to understand complex relationships.
The training pipeline first constructs the negative graph to ensure a balanced training dataset by mimicking real-world scenarios where users do not interact with all items. We use the Adam optimizer with weighted decay to optimize margin-based loss, where the positive edge scores are encouraged to exceed negative edge scores by at least a margin of 1, penalizing cases where this condition is not met.
During evaluation, the ZiGNN’s effectiveness was tested using node embeddings derived from the trained model. These embeddings were normalized to enable cosine similarity-based retrieval, crucial for recommendation tasks. For specific canonical edge types (e.g., ’user’, views, ’city’), embeddings were computed separately for source (e.g., user) and destination nodes (e.g., city). The recommendation performance of ZiGNN is measured using nDCG@K, which evaluates the ranking quality of recommendations and prioritizes highly relevant items. Metrics like nDCG@K was used to evaluate recommendation quality, providing quantitative insights.
Z-REx explains the recommendations generated by ZiGNN, and for this experimentation, we focus on user-to-city views relationships, but this works for any relationships. First, we identify an important subset of features by measuring the change in nDCG@K score when individual features are zeroed out. Then, selecting those that produce a negative impact on nDCG, meaning those features were important for predicting relevant recommendations. To align with the recommendation explainability task, we redefine the traditional Fidelity metric (yuan2021explainability)—originally based on change in prediction confidence—as the change in nDCG@K score due to perturbations. The explanation performance of Z-REx is then evaluated based on the drop in nDCG@K score and the change in cosine similarity ($\Delta\text{sim}$) between user and city embeddings caused by structural perturbations and by only using the important features.
6.2. Dataset
The dataset was collected over a three-day period from the state of Washington in the USA. The dataset statistics are shown in Table 4, and more details in § A.2. The dataset is split into training, validation, and test sets, ensuring that the test set contains previously unseen pairs, which are evaluated on the following day to avoid any data leakage.
Focusing on the training data some interesting and some expected trends are seen. Expected trends are seen such as view events are the most popular event, followed by save and then toured because users tend to view multiple items before saving them and finally touring the item. The average interaction of the user to listing for view event for the 25th quantile is 1.0 and the 75th quantile is 8.0, the save event for the 25th quantile is 2.7 and the 75th quantile is 3.0, and the tour event for the 25th quantile is 3.2 and the 75th quantile is 4.0. Therefore, from the average interaction of different events statistics tells us that the graph generated from the dataset will contain more sparse connections than dense connections.
Figure 5. Performance of ZiGNN against different baselines.
6.3. Recommendation Accuracy of ZiGNN
To investigate RQ1, we compared the performance between the ZiGNN and the baseline recommendation models to reveal ZiGNN consistently and substantially outperforms the rest as seen in Figure 5. ZiGNN consistently surpasses the histogram model across all values of K, with the performance gap widening as K increases. Notably, at $K=10$, the ZiGNN achieves an nDCG score of 0.019, a 52% improvement over the histogram model’s 0.0125, demonstrating ZiGNN’s ability to produce more accurate and relevant recommendations as the recommendation set grows.
This trend is further emphasized by the steep growth curve of the GNN’s performance compared to the relatively linear increase observed for the histogram model, showcasing the ZiGNN’s adaptability. Interestingly, even at $K=1$, where only the top recommendation matters, the ZiGNN outperforms the histogram model by a remarkable 200% (0.0075 vs. 0.0025), highlighting its capacity to prioritize the most relevant results effectively. These findings suggest that the ZiGNN’s structural representation learning captures nuanced relationships in the data that the histogram model, with its simpler statistical approach, fails to address.
Figure 6. Impact of zeroing out features to find important features.
Figure 7. Impact of sequentially adding important features on nDCG.
6.4. Explanation Quality of Z-REx
For investigating RQ2, we first consider the feature perturbation quality and show that we can identify the top ten important city features for the recommendation by measuring the difference between the original nDCG and the perturbed nDCG when a feature is zeroed out. When the difference is positive, the feature was important for predicting the recommendation, as shown in Figure 6, and features with negative differences mean they were confusing the model (more details in § A.3). In particular, these important features can be grouped: (1) geographic - latitude (top right and bottom left) and longitude (bottom left), (2) boolean - view, spa, and vacant land, and (3) numeric - year, population count, days on the market, and square feet. As illustrated in Figure 7, using only the important features yield better performance than using all the features, demonstrating that Z-REx effectively identifies the important features needed for user-city recommendation. These top features will be the only features used for measuring the structural perturbation quality.
The important city features align with domain knowledge: a city is typically characterized by its geographic location and aggregate properties (e.g., population count, the average year of houses, and average days on the market), which are different from the features usually used to characterize a listing (e.g., the number of bathrooms and bedrooms). Aggregate properties offer a holistic snapshot of a city’s dynamics and freshness. Interestingly, cities with a younger infrastructure experience rapid growth, indicating a strong interplay between user preference and given recommendations. Moreover, as shown in Figure 9, we notice the unimportant features that confuse ZiGNN are listing features—removing them leads to improved performance.
nDCG (decr.) 94% 81% (-13%) 21% (-73%) 47% (-47%)
cosine similarity -0.10 -0.07 (-0.03) -0.02 (-0.08) -0.04 (-0.06)
Table 1. Quantitative evaluation of Z-REx against GNN explainers.
6.5. SOTA GNN Explainers vs. Z-REx
We compared Z-REx against a specialized GNN-based recommendation explainer PaGE-Link (zhang2023page) and two general GNN explainers, GNNExplainer (ying2019gnnexplainer) and SubgraphX (yuan2021explainability) in Table 1 to answer RQ3. Z-REx performed the best, followed by PaGE-Link and general purpose GNN explainers performed the worst which is expected since Z-REx and PaGE-Link have been specifically designed for LP tasks. Compared to GNNExplainer, Z-REx demonstrates a -73% more reduction in nDCG and -0.08 more cosine similarity increase, indicating that it more effectively identifies the features and subgraphs the GNN relies on to make recommendations. Similarly, against SubgraphX, Z-REx shows greater degradation in both metrics. Compared to PaGE-Link, Z-REx the comparison performance are relatively closer, showing PaGE-Link is better than both SubgraphX and GNNExplainer. Z-REx has 13% more decrease in nDCG is because Z-REx considers the entire graph and all the available features as context and since PaGE-Link is limited by the ego-graph size, it is unable to consider some of the relevant features that are important to the GNN. This limitation also impacts PaGE-Link in identifying relevant subgraph structures which is reflected by -0.03 less decrease is cosine similarity. These results highlight Z-REx’s distinct advantages over existing explainers in altering model behaviors and disrupting reliance on original explanations.
Figure 8. Z-REx’s explanation of a recommended city #1.
6.6. Case Study
In the case study, we selected a real user and based on their interactions generated city recommendations and explanations for those recommendations. To simplify the visualization, we compress the graph to only the specified user (yellow node) and their co-clicked cities (blue nodes). By co-clicked cities, we mean cities the current user clicked listings in as well as cities other users clicked listings in where they share at least one common city with the current user. A blue edge will exist between the co-clicked city nodes, the recommended city is the green node, and the red edges are the ones we have identified as important. To enhance visualization, we present a simplified graph (Figure 8) focusing on the target user (yellow node), co-clicked cities (blue nodes), and the recommended city (green node). A co-clicked city indicates that the user interacted with listings in both cities, implying a potential relationship. Red edges represent the connections identified by Z-REx as most influential in the recommendation.
7. Conclusion
We introduced Z-REx, which aims to bridge a critical gap in explainable ML for recommendation systems by providing human-interpretable insights into GNN-based recommenders to increase user confidence and customer engagement. Specifically, we provide whole graph instance-level explanations for GNN models in the context of the heterogeneous link prediction (LP) tasks. Z-REx offers contextually relevant and human-interpretable explanations through structural and attribute perturbation. We also introduce ZiGNN, a GNN-based recommendation system tailored for real-world real-estate heterogeneous datasets. We use Z-REx to explain the recommendations of ZiGNN to show its superior explanation performance compared to the SOTA GNN-explainers by providing qualitative insights and quantitative performance.
@@ 7. Conclusion
[heading=none]
Appendix A Appendix
Src Node Type Dst Node Type Edge Type
User Listing views, saves, and tours
User City searched in
City Listing contains
Table 2. Description of relationships in the heterogeneous graph.
@@ Appendix A Appendix
A.1. Node and Edge Details
@@
Date City Listing User views saves tours contains
Training 5/17-5/20 449 55k 393k 789k 169k 234k 55k
Testing 5/27-5/30 452 53k 405k 773k 138k 197k 53k
Evaluation 5/31 448 45k 203k 714k 41k 56k 45k
Table 4. Dataset summary for training, testing, and evaluation data.
A.2. Dataset Statistics
The dataset statistics as shown in Table 4 shows that the datset is divided into three segments—training (May 17–20), testing (May 27–30), and evaluation (May 31)—providing a temporal snapshot of various metrics. The training dataset, spanning May 17th to May 20th, includes 55k listings viewed by 393k users across 449 cities, resulting in 789k views, 169k saves, and 234k tours. The contains relationship matches the number of listings since each listing must belong to a city. The testing dataset exhibits similar characteristics from May 27th to May 30th with slightly different values (53k listings and 405k users). Notably, the evaluation dataset on May 31st, while covering a similar number of cities (448), shows a smaller number of listings (45k), users (203k), saves (41k), and tours (56k) in comparison to training and testing datasets since there is a decrease in user engagement during the evaluation period which is just one day after the testing period. The views are the largest interactions since users view the listing the most, and only a few of them convert to saves, and fewer convert to tours.
Table 3 outlines the attribute schema for the different node types used in our model. The User node is characterized by a numeric session identifier. The Listing node includes a diverse set of features: numeric attributes (e.g., bedrooms, bathrooms, price, etc.), boolean attributes (e.g., waterfront, heating, etc.), and detailed geographic attributes (latitude and longitude for the property boundaries). Meanwhile, the City node aggregates listing data by averaging numeric, boolean, and geographic attributes and providing a population count. This structured representation supports our approach to integrating heterogeneous data for improved node representation learning in the ML framework.

### v2

@@ A.1. ZiGNN Architecture
During training, both positive and negative edge scores are calculated for each edge type by applying the dot-product predictor to pairs of node embeddings, enabling the model to distinguish between observed and unobserved relationships. We use the Adam optimizer with weighted decay to optimize margin-based loss, where the positive edge scores are encouraged to exceed negative edge scores by at least a margin of 1, penalizing cases where this condition is not met. The modular design supports inference by directly returning node embeddings when negative edges are not provided. This architecture ensures the model can efficiently learn and generalize from multi-relational graph data for edge classification tasks.
Algorithm 2 Z‑REx
1:target user $u$, target city $c_{t}$, heterogeneous interaction graph $G_{\text{hetero}}$, trained GNN ZiGNN, rank cut‑off $K$, city‑feature $F_{\text{all}}$, hop distance $k$
2:function Z‑REx($u,c_{t},F_{\text{all}},k$)
@@
[…]
6:end function
7:function FeaturePerturb($u,c_{t},F_{\text{all}}$)
8:     $baseline\_ndcg\leftarrow\textsc{NDCG}(u)$triangleright original score
[…]
10:     for all $f\in F_{\text{all}}$ do
11:         $\bar{x}_{c_{t}}\leftarrow\textsc{ZeroOutFeature}(c_{t},f)$
12:         $\Delta ndcg\leftarrow\textsc{NDCG}(u,\bar{x}_{c_{t}})-baseline\_ndcg$triangleright Eq.(3)
[…]
14:     end for
15:     Sort(ranked_features) by $\Delta ndcg$ (descending)
16:     return ranked_features
17:end function
18:function StructurePerturb($u,c_{t},k$)
19:     $G_{h}\leftarrow\textsc{CollapseToUserCityGraph}(G_{\text{hetero}},c_{t})$
20:     $G_{k}\leftarrow\textsc{KHopSubgraph}(G_{h},u,k)$
21:     AddCo-ClickEdges($G_{k}$)
22:     $base\_sim\leftarrow\text{cos}(\textsf{ZiGNN.embed}(u),\textsf{ZiGNN.embed}(c_{t}))$
[…]
24:     for all edge $e$ incident to $u$ or $c_{t}$ in $G_{k}$ do
25:         temporarily remove $e$
26:         $new\_sim\leftarrow\text{cos}(\textsf{ZiGNN.embed}(u),\textsf{ZiGNN.embed}(c_{t}))$
27:         $\Delta sim\leftarrow new\_sim-base\_sim$triangleright Eq.(4)
[…]
29:         restore $e$
30:     end for
31:     Sort(ranked_edges) by $|\Delta sim|$ (descending)
32:     return ranked_edges
33:end function
Algorithm 3 Generate Heterogeneous Interaction Graph
1:CSVs (i.e., $events\_file$, $regions\_file$), normalization method $norm\_method$
2:function Preprocess($events\_file,regions\_file$)
3:     $events\leftarrow\textsc{ReadCSV}(events\_file)$
4:     $regions\leftarrow\textsc{ReadCSV}(regions\_file)$
5:     $U_{L}\leftarrow\textsc{BuildUserListingEdges}(events)$
6:     $regions^{\prime}\leftarrow regions[id\in U_{L}.listing\_id]$
7:     $U_{C}\leftarrow\textsc{BuildUserCityEdges}(U_{L},regions^{\prime})$
8:     $C_{L}\leftarrow\textsc{BuildCityListingEdges}(regions^{\prime})$
9:     $users\leftarrow\textsc{Unique}(events.user\_id,Xavier\_Init)$
10:     $listings\leftarrow\textsc{CleanNumeric}(regions^{\prime},norm\_method)$
11:     $cities\leftarrow\textsc{AggregateCity}(regions^{\prime},norm\_method)$
12:     $G_{hetero}\leftarrow\textsc{dgl.heterograph}(U_{L},U_{C},C_{L})$
13:     return $G_{hetero}$
14:end function
A.2. Z-REx Algorithm
Algorithm 2 describes the Z-REx algorithm and Algorithm 3 describes the algorithm to generate the heterogeneous interaction graph, $G_{hetero}$ from clickstream events. Algorithm 3 converting raw interaction and regional metadata CSVs into attribute-normalized heterogeneous graph, $G_{hetero}$, that connects the users, listings, and cities through user-listing, user-city, and city-listing edges.
Using $G_{hetero}$, Algorithm 2 delivers fine-grained model transparency: for a given user-city recommendation, it (1) ranks city-level features by zeroing them out and measuring the resulting $\Delta nDCG@K$, and (2) ranks graph edges in the $K$-hop neighbourhood by perturbing edge and observing the change in the cosine similarity between the ZiGNN embeddings of the user and target city. The joint output—two ordered lists of high-impact features and structural subgraphs, exposes the concrete evidence that the GNN relied upon.
@@ A.2. Z-REx Algorithm
A.3. Node and Edge Details
@@
Table 4 outlines the attribute schema for the different node types used in our model. The User node is characterized by a numeric session identifier. The Listing node includes a diverse set of features: numeric attributes (e.g., bedrooms, bathrooms, price, etc.), boolean attributes (e.g., waterfront, heating, etc.), and detailed geographic attributes (latitude and longitude for the property boundaries). Meanwhile, the city node aggregates listing data by averaging numeric, boolean, and geographic attributes and adding population count.
Date City Listing User
3-days Training 5/17-5/20 449 55k 393k
30-days Training 4/20-5/20 456 97k 1.3M
Testing 5/27-5/30 452 53k 405k
Evaluation 5/31 448 45k 203k
Table 5. Number of entities (e.g., nodes) in different datasets.
Date views saves tours contains
3-days Training 5/17-5/20 789k 169k 234k 55k
30-days Training 4/20-5/20 26M 2.2M 1.6M 97k
Testing 5/27-5/30 773k 138k 197k 53k
Evaluation 5/31 714k 41k 56k 45k
Table 6. Number of relationships (e.g., edges) in different datasets.
views saves tours
3-days
mean 7.97 2.74 3.24
25th quantile 1.00 1.00 1.00
median 3.00 2.00 2.00
75th quantile 8.00 3.00 4.00
30-days
mean 21.90 5.65 6.72
25th quantile 2.00 1.00 1.00
median 5.00 2.00 3.00
75th quantile 18.00 6.00 7.00
Table 7. Average number of user-city interactions.
A.4. Dataset Statistics
The dataset statistics as shown in Table 5 and Table 6 shows that the dataset is divided into four segments: 3-day training (May 17-20), month-long training (April 20 - May 20), testing (May 27-30), and evaluation (May 31)—providing a temporal snapshot of various metrics. The 3-day training dataset, spanning May 17th to May 20th, includes 55k listings viewed by 393k users across 449 cities, resulting in 789k views, 169k saves, and 234k tours. The contains relationship matches the number of listings since each listing must belong to a city. The testing dataset exhibits similar characteristics from May 27th to May 30th with slightly different values (53k listings and 405k users). Notably, the evaluation dataset on May 31st, while covering a similar number of cities (448), shows a smaller number of listings (45k), users (203k), saves (41k), and tours (56k) in comparison to training and testing datasets since there is a decrease in user engagement during the evaluation period which is just one day after the testing period. The views are the largest interactions since users view the listing the most, and only a few of them convert to saves, and fewer convert to tours.
As illustrated in Table 7, user engagement with listings scales markedly with the observation window. Over the 7-day horizon, users perform a mean of 7.97 views, 2.74 saves, and 3.24 tour requests per listing, yet the corresponding medians (3, 2, 2) reveal a highly right-skewed distribution in which a small subset of users interactions form the majority of the engagement dataset. Extending the window to 30 days amplifies this effect: mean interactions roughly triple for views and more than double for both saves and tours, while the medians increase more modestly (to 5, 2, 3), indicating that additional time primarily benefits heavier-engagement cohorts. The widening gap between the 25th and 75th percentiles across all metrics further shows the importance of the growing dispersion, suggesting that long-term recommendation strategies should explicitly account for this heterogeneity rather than relying on aggregate averages alone.
@@ A.5. Unimportant Features
Table 8. ZiGNN Runtime performance on different dataset windows.
Scenario # of Edges Time (s)
3‑days
Training per epoch 1.2M 0.88
Evaluation per user 248k 5.14
30‑days
Training per epoch 29.8M 10.89
Evaluation per user 248k 5.04
A.6. Runtime Analysis: 3 vs 30 Day Dataset
Table 8 shows that enlarging the interaction window from 3 days (1.2M edges) to 30 days (29.8M edges) raises the cost of one training epoch from 0.88s to only 10.89s. Despite a 25 times increase in graph size, the runtime grows by just one order of magnitude, confirming that our training pipeline scales sub‑linearly with temporal coverage. Inference remains virtually size‑independent: generating an explanation for a single user takes 5.14s on the 3‑day graph versus 5.04s on the 30‑day graph. Together, these results demonstrate that ZiGNN delivers timely and accurate explanations at production scale.